OpenStudy (anonymous):

Evaluate: 10 − 64 ÷ 8 + 42

OpenStudy (unklerhaukus):

10 − 64 ÷ 8 + 42 = 10 − (64 ÷ 8) + 42

OpenStudy (kuoministers):

Its simple use BODMAS (its the order in which you do this question). It goes Brackets pOwers Division Multiplication Addition then Subtraction. So you would do the division first. 10 - (64 / 8) + 42. 10 - 8 + 42 Then you can do the rest

OpenStudy (anonymous):

=10 - (64 / 8) + 42 =10 - 8 + 42 =2 + 42 = 44

OpenStudy (kuoministers):

well.. yes....